### v2

# Discovering Effective Policies for Land-Use Planning

Abstract
How areas of land are allocated for different uses, such as forests, urban, and agriculture, has a large effect on carbon balance, and therefore climate change. Based on available historical data on changes in land use and a simulation of carbon emissions/absorption, a surrogate model can be learned that makes it possible to evaluate the different options available to decision-makers efficiently. An evolutionary search process can then be used to discover effective land-use policies for specific locations. Such a system was built on the Project Resilience platform (ITU, 2023) and evaluated with the Land-Use Harmonization dataset (Hurtt et al., 2020) and the BLUE simulator (Hansis et al., 2015). It generates Pareto fronts that trade off carbon impact and amount of change customized to different locations, thus providing a potentially useful tool for land-use planning.
1 Introduction
One of the main factors contributing to climate change is how much land area is allocated for different uses (Friedlingstein et al., 2022). Forests in general absorb more carbon than e.g. crops and ranges, yet such uses are essential for the economy. Land-use patterns must therefore be planned to maximize carbon absorption while maintaining economic viability.
An approach for land-use optimization was developed as part of Project Resilience, a non-profit project hosted by the ITU agency of the United Nations (ITU, 2023). The goal is to provide decision-makers with a tool to know how their land-use choices affect CO2 fluxes in the long term, and make suggestions for optimizing these choices.
More specifically, the tool is designed to answer three questions: (1) For a geographical grid cell, identified by its latitude and longitude, what changes to the land usage can be made to reduce CO2 emissions? (2) What will be the long-term CO2 impact of changing land usage in a particular way? (3) What are the optimal land-use choices that can be made with minimal cost and maximal effect?
The approach is based on the Evolutionary Surrogate-assisted Prescription method (Francon et al., 2020). The idea is to first utilize historical data to learn a surrogate model on how land-use decisions in different contexts affect carbon emissions. Then, this model is used to evaluate candidates in an evolutionary search process for good land-use change policies.
As a result, a Pareto front is generated of solutions that trade off reduction in carbon emissions and the amount of change in land use. Each point in the Pareto front represents an optimal policy for that tradeoff. To make the results trustworthy, the tool allows the decision-maker to explore modifications to these policies, and see the expected effect. In the future, it should also be possible to evaluate confidence of the predictions, evolve rule sets to make the policies explainable, and utilize ensembling and further objectives and preferences to make them more accurate. Thus, the tool harnesses several techniques in machine learning to provide a practical tool for decision-makers in optimizing land-use decisions.
An interactive demo of the system is at https://landuse.evolution.ml.
2 Background
2.1 Evolutionary Surrogate-assisted Prescription
Figure 1: The ESP Decision Optimization Method. A predictor is trained with historical data on how given actions in given contexts led to specific outcomes. It is then used as a surrogate in order to evolve prescriptors, i.e. neural networks that implement decision policies resulting in the best possible outcomes.
Evolutionary Surrogate-assisted Prescription (ESP; Francon et al., 2020) is an approach for optimizing decision-making in a variety of domains (Figure 1). The main idea is that a decision policy can be represented as a neural network, or a set of rules, and a good policy can be discovered through population-based search, i.e. using evolutionary computation techniques. However, each candidate must be evaluated, which is difficult to do in many real-world applications. Therefore, a surrogate model of the world is learned from historical data, predicting how good the resulting outcomes are for each decision in each context.
More formally, given a set of possible contexts $\mathbb{C}$ and possible actions $\mathbb{A}$, a decision policy $D$ returns a set of actions $A$ to be performed in each context $C$:
$D(C)=A\;,$ (1)
where $C\in\mathbb{C}$ and $A\in\mathbb{A}$. For each such $(C,A)$ pair there is a set of outcomes $O(C,A)$, and the Predictor $P_{d}$ is defined as
$P_{d}(C,A)=O,$ (2)
and the Prescriptor $P_{s}$ implements the decision policy as
$P_{s}(C)=A\;,$ (3)
such that $\sum_{i,j}O_{j}(C_{i},A_{i})$ over all possible contexts $i$ and outcome dimensions $j$ is maximized (assuming they improve with increase). It thus approximates the optimal decision policy for the problem. The predictor can be learned from historical data on $CAO$ triples. In contrast, the optimal actions $A$ for each context $C$ are not known, and must therefore be found through search.
ESP was first evaluated in reinforcement learning tasks such as function approximation, cart-pole control, and the flappybird game, and found to discover significantly better solutions, find them faster, and with lower regret than standard approaches such as direct evolution, DQN, and PPO. Its power comes from automatic regularization and automatic curricular learning: When predictors are learned at the same time as the prescriptors, each prescriptor is evaluated against multiple predictors; as the predictors improve, they provide more refined evaluations. ESP is thus a powerful mechanism for learning policies in time-varying domains.
2.2 Project Resilience
A major application of ESP to decision-making was developed for optimizing strategies for non-pharmaceutical interventions (NPIs) in the COVID-19 pandemic (Miikkulainen et al., 2021). Using case and NPI data from around the world, collected in 2020-2022 by Oxford University’s Covid-19 project (Hale et al., 2020), a neural network model was developed to predict the course of the pandemic; this network was then used as a surrogate in evolving another neural network to prescribe NPIs. It discovered different strategies at different phases of the pandemic, such as focusing on schools and workplaces early on, alternating policies over time, and focusing on public information and masking later. This application demonstrated the power of ESP on discovering good tradeoffs for complex real-world decision-making tasks.
Encouraged by this application, the XPRIZE Pandemic Response Challenge (XPRIZE, 2023) was developed in 2020-2021 for NPI optimization using the same predictor+prescriptor approach. The competitors used several different machine learning or other techniques, and drew significant attention to AI-supported decision-making. The results were effective and in some cases deployed by decision-makers. An important conclusion was that it was possible to convene a group of experts – data scientists, epidemiologists, and public health officials, to build a useful set of tools to advise us on how to cope with and plan around a health disruption to society.
This experience with the XPRIZE Pandemic Response Challenge then led to Project Resilience (ITU, 2023). The XPRIZE competition serves as a blueprint for a collaborative effort to tackle global problems. The goal of Project Resilience is to build a public AI utility where a global community of innovators and thought leaders can enhance and utilize a collection of data and AI approaches to help with better preparedness, intervention, and response to environmental, health, information, social equity, and similarly scoped problems. The project is hosted by the ITU agency of the United Nations, under its Global Initiative for AI and Data Commons, and contributes to its general efforts towards meeting the Sustainable Development Goals (SDGs).
Staffed by volunteers from around the world, Project Resilience was started in 2022 with the design of a minimum viable product (MVP) platform. The goal is to
  • Develop an architecture to pull input and output data hosted by third parties;
  • Build an API for third parties to submit models.
  • Develop code to compare both predictors and prescriptors in third-party models and produce a set of performance metrics;
  • Build a portal to visualize assessment of predictors and prescriptors to include generations of key performance indicators (KPIs) and comparison across models;
  • Develop an ensemble model for predictors and prescriptors; and
  • Provide a user interface to decision-makers to help them in their decision-making process.
Land-use optimization is the first application of this MVP, as described below.
3 Land-Use Optimization Task
The data sources are described first, followed by how they are used as the context, action, and outcome variables. The elements of the approach include the predictor and prescriptor models, a method for modeling uncertainty, and ensembling the results.
3.1 Data
The data for carbon emissions (Emissions resulting from Land-Use Change, ELUC) originate from a high-fidelity simulation called Bookkeeping of Land-Use Emissions (BLUE; Hansis et al., 2015). BLUE is designed to estimate the long-term CO2 impact of committed land use. “Committed emissions” means all the emissions that are caused by a land-use change event are attributed to the year of the event. BLUE is a bookkeeping model that attributes carbon fluxes to land-use activities. While in principle a simulator can be used as the surrogate model for ESP, in practice the simulations are too expensive to carry out on demand during the search for good policies. Therefore, the team in charge of the BLUE model performed a number of simulations covering a comprehensive set of situations for 1850-2022, resulting in a dataset that could be used to train an efficient surrogate model.
The Land-Use Change (LUC) data is provided by the Land-Use Harmonization project (LUH2; Hurtt et al., 2020). A land-use harmonization strategy estimates the fractional land-use patterns, underlying land-use transitions, and key agricultural management information, annually for the time period 850-2100 at 0.25 x 0.25 degree resolution.
The most recent version of the land-use data (2022) consists of separate layers for states, transitions, and management. These files are based on the new HYDE3.3 historical database, as well as the 2021 FAO wood harvest data, for all years 850-2022. It consists of the following land-use types:
Primary: Vegetation that is untouched by humans
  • primf: Primary forest
  • primn: Primary nonforest vegetation
Secondary: Vegetation that has been touched by humans
  • secdf: Secondary forest
  • secdn: Secondary nonforest vegetation
Urban
  • urban: Urban areas
Crop
  • c3ann: Annual C3 crops (e.g. wheat)
  • c4ann: Annual C4 crops (e.g. maize)
  • c3per: Perennial C3 crops (e.g. banana)
  • c4per: Perennial C4 crops (e.g. sugarcane)
  • c3nfx: Nitrogen fixing C3 crops (e.g. soybean)
Pasture
  • pastr: Managed pasture land
  • range: Natural grassland / savannah / desert / etc.
Together the ELUC and LUC datasets form the basis for constructing the context, action, and outcome variables, as will be described next.
3.2 Decision-making Problem
The modeling approach aims to understand the domain in two ways: (1) In a particular situation, what are the outcomes of the decision maker’s actions? (2) What are the actions that result in the best outcomes, i.e. the lowest carbon emission (or highest capture) with the smallest change? The data is thus organized into context, action, and outcome variables.
Context describes the problem the decision maker is facing, i.e. a particular grid cell, a point in time when the decision has to be made, and the usage of the land at that point. More specifically it consists of:
  • Latitude and Longitude, representing the cell on the grid.
  • Area, representing the surface of the cell. Cells close to the equator have a bigger area than cells close to the poles.
  • Year, useful to capture historical decisions: The same cell may have been through many land-use changes over the years.
  • Land usage, representing the percentage of the land used by each land type above. In addition to those listed above, there is a ‘nonland’ type that represents the percentage of the cell that is not land (e.g. typically sea, lake, etc.). The land usage types sum up to 100%.
Actions represent the choices the decision-maker faces. How can they change the land? In the study of this paper, these decisions are limited in two ways:
First, decision-makers cannot affect primary land. The idea is that it is always better to preserve primary vegetation; destroying it is not an option given to the system. Technically, it is not possible to re-plant primary vegetation. Once destroyed, it is destroyed forever. If re-planted, it would become secondary vegetation.
Second, decision-makers cannot affect urban areas. The needs of urban areas are dictated by other imperatives, and optimized by other decision makers. Therefore, the system cannot recommend that a city should be destroyed, or expanded.
Outcomes consist of two conflicting variables. The primary variable is ELUC, i.e. emissions from land-use change. It consists of all CO2 emissions attributed to the change, in metric tons of carbon per hectare (tC/ha), obtained from the BLUE simulation. A positive number means carbon is emitted, a negative number means carbon is captured. The secondary variable is the cost of the change, represented by the percentage of land that was changed. This variable is calculated directly from the actions.
There is a trade-off between these two objectives: It is easy to reduce emissions by changing most of the land, but that would come at a huge cost. Therefore, decision-makers have to minimize ELUC while minimizing land change at the same time. Consequently, the result is not a single recommendation, but a Pareto front where each point represents the best implementation of each tradeoff given a balance between the two outcomes.
4 Models
The system consists of the predictor, trained with supervised learning on the historical data, and the prescriptor, trained through evolution.
Prediction:
Given the context and actions that were performed, the predictive model estimates the outcomes. In this case, since the cost outcome can be calculated directly, only the ELUC is predicted by the model. That is, given the land usage of a specific location, and the changes that were made during a specific year, the model predicts the CO2 long-term emissions directly caused by these changes.
Any predictive model can be used in this task, including a neural network, random forest, or linear regression. As usual, the model is fit to the existing historical data and evaluated with left-out data.
Prescription:
Given context, the prescriptive model suggests actions that optimize the outcomes. The model has to do this for all possible contexts, and therefore it represents an entire strategy for optimal land use. The strategy can be implemented in various ways, including decision trees, sets of rules, or neural networks. The approach in this paper is based on neural networks.
The optimal actions are not known, but the performance of each candidate strategy can be measured (using the predictive model), therefore the prescriptive model needs to be learned using search techniques. Standard reinforcement learning methods such as PPO and DQN are possible; the experiments in this paper use evolutionary optimization, i.e. conventional neuroevolution (Stanley et al., 2019). As in prior applications of ESP (Francon et al., 2020; Miikkulainen et al., 2021), the network has a fixed architecture of two fully connected layers; its weights are concatenated into a vector and evolved through crossover and mutation.
5 Experiments
The main (early) results in the project so far, on different ways to create good predictors, on evolving effective prescriptors, and on an interactive interface for analyzing the results, are described in this section, with the next steps outlined in Section 6.
5.1 Prediction
Model Train Time(s) W. Europe South America United States Global
LinReg (EU) 0.1236 0.0331 0.2570 0.1720 0.2204
LinReg (SA) 0.8281 0.1422 0.1549 0.0648 0.1110
LinReg (US) 0.5652 0.1402 0.1467 0.0345 0.0720
LinReg (Global) 11.1449 0.1410 0.1520 0.0366 0.0723
RF (EU) 43.2344 0.0314 0.2373 0.1156 0.2232
RF (SA) 338.2790 0.1722 0.0715 0.0462 0.1095
RF (US) 115.4622 0.1559 0.1874 0.0200 0.1120
RF (Global) 255.3720 0.1007 0.0870 0.0257 0.0558
Neural Net (EU) 78.7409 0.0247 0.5510 0.3313 0.3493
Neural Net (SA) 1208.0376 0.7950 0.0936 0.1866 0.2195
Neural Net (US) 312.3967 0.3348 0.2164 0.0196 0.1319
Neural Net (Global) 186.2248 0.3298 0.2182 0.1309 0.0530
Table 1: Mean absolute error of the models trained on each region, evaluated on that region as well as all other regions. The Neural Net models are more accurate than the RF models (with the exception of SA), which are more accurate than LinReg, and the specialized models are more accurate than the global model. The Neural Net models both capture nonlinearities and extrapolate well in extreme cases; therefore, the Global Neural Net model was used to evolve the prescriptors.
In preliminary experiments, it turned out to be difficult to achieve high accuracy globally in a single model. Therefore, separate models were trained on different subsets of countries: Western Europe, South America, and the United States. The Western Europe subset contains the UK, France, Germany, the Netherlands, Belgium, Switzerland, and Ireland. The South American subset includes Brazil, Bolivia, Paraguay, Peru, Ecuador, Colombia, Venezuela, Guyana, Suriname, Uruguay, Argentina, and Chile. These countries were chosen as a representative subset to make the experiments computationally less expensive.
Upon examining the dataset, a linear relationship was found between ||||ELUC|||| and change with an R-squared value of 0.92. Therefore, a linear regression (LinReg) was fit to the data as the first predictive model. In preliminary experiments, the LinReg model was found to perform the best when current land use and cell area were removed from the inputs. Thus, the differences in land use were used as the only input features in the experiments. Data from years [1851-2011] was used to train the LinReg models.
Second, a random forest model (RF) was trained on the same data, with the default scikit-learn parameters and the full set of input features. To make training time feasible, the models for Western Europe, South America, and the United States were trained on data from the years [1982-2011] and the global models were trained with data from [2010-2011]. Preliminary experiments showed that constraining the training set to more modern years did not significantly decrease performance.
Third, a neural network (neural net) was trained for the same task. For Western Europe, South America, and the United States, it was trained on the same data as LinReg; For global, the neural net was trained on a random 10% of the global data, due to training time limitations. The neural nets were fully connected, composed of an input layer, a single hidden layer of size 4096, and an output layer that took as input the concatenation of the input layer and the hidden layer in order to take advantage of the observed linear relationships more easily (Cheng et al., 2016).
After training, all models were tested with data from [2012-2021]. As shown in Table 1, the LinReg models performed consistently the worst, suggesting that the problem includes significant nonlinear dependencies after all. RF performed significantly better; however, RF models do not extrapolate well beyond the training examples, which turns out to be important when evolving prescriptors for extreme land-use changes. In contrast, neural nets both capture nonlinearities and extrapolate well, and turned out to be best models overall. In each region the neural network trained on that region performed the best, except in SA, where it was close. The LinReg and neural network have many fewer parameters than RF (taking 4KB and 440KB on disk respectively vs. 2GB for RF), and LinReg trains the fastest.
Figure 2: Visualizing the differences in model behavior. Predictions are created for the global models using synthetic data that consists of changing 100% of land type A (row) to 100% of type B (column). The models generally agree on the sign of ELUC, suggesting that the results are reliable. The RF model is not able to extrapolate to extreme values, resulting in low predictions; LinReg and Neural Net are similar but differ numerically, presumably due to the differences between linear and nonlinear predictions.
In order to visualize the behavior of the predictor models, they were queried by every possible conversion from 100% type A to 100% type B (the types primf, primn, and urban were excluded from B and c4per from A and B for reasons outlined in Section 5.2). Figure 2 shows the result for the global models. The main result is that all models generally rank crops $<$ pasture $<$ forest in terms of reducing ELUC, and generally agree on the sign of ELUC. This consistency suggests that the results are reliable. Second, while LinReg and Neural Net make similar predictions overall, their numerical values differ in many places, presumably due to the nonlinearities. Third, despite RF’s generally high accuracy in known scenarios, it is not able to extrapolate well to extreme changes in land usage, resulting in low-magnitude ELUC values.
Thus, since the neural network achieves better performance than the RF in most cases, extrapolates well to extremes, and can learn nonlinear relationships in the data, the global neural network was used as the surrogate to evolve the prescriptors.
5.2 Prescriptions
In the experiments in this paper, the prescriptors were fully connected neural networks with two layers of weights, implemented with Keras/Tensorflow. The topology was fixed and only the weights were evolved. The input consisted of 13 float values, i.e. the 12 land-usage percentages and the area of the geographical cell. The hidden layer contained 16 units with tanh activation.
During evolution, prescriptor candidates were evaluated on a random subset of 1% of the global dataset from the years [1851-2011] with the global neural net predictor. Prescriptor output was a single vector of size eight (the 12 land usage percentages minus primf, primn, and urban, which were considered fixed, and c4per which was found to occur too rarely in the dataset to predict reliably), and a softmax activation. These eight outputs specify the prescriptor’s suggestion for the percentage of land usage of each type. The eight outputs were combined with the four fixed land-usage values and the cell area and passed to the Predictor model to get the predicted ‘ELUC’ metric. This metric is aggregated across the evaluation data set to form the performance outcome of the suggestion. The difference between the eight outputs and the corresponding inputs is used to compute the percentage of land-use change, i.e. the cost outcome of the suggestion.
The aggregated ELUC and the change outcomes were minimized with an evolutionary search on the neural net weights. Conventional neuroevolution with crossover and mutation operators on a vector of network weights was combined with NSGA-II (Deb et al., 2000) to account for multiobjectivity. To speed up evolution, two prescriptor models were trained with backpropagation: One to prescribe no changes and the other to prescribe as much secdf as possible (which was observed to reduce ELUC the most); these two models were injected into the initial population. The rest of the population was created randomly through an orthogonal initialization of weights in each layer with a mean of 0 and a standard deviation of 1 (Saxe et al., 2014). Evolution was run for 100 generations with the following parameters:
  • nb_elites: 10
  • mutation_type: gaussian_noise_percentage
  • mutation_factor: 0.2
  • population_size: 100
  • parent_selection: tournament
  • initialization_range: 1
  • mutation_probability: 0.2
  • remove_population_pct: 0.8
  • initialization_distribution: orthogonal
Figure 3: Evolution of prescriptors with the global neural network predictor. (Left) The Pareto front moves towards the lower left corner over evolution, finding better implementations for the different tradeoffs of the ELUC and change objectives. (Right) Each prescriptor evaluated during evolution is shown as a dot, demonstrating a wide variety of solutions and tradeoffs. The final Pareto front is shown as red dots in both figures, constituting a set of solutions from which the decision-maker can choose a preferred one.
Figure 3 demonstrates the progress of evolution towards increasingly better prescriptors, i.e. those that represent better implementations of each tradeoff of the ELUC and change objectives. They represent a wide variety of tradeoffs, and a clear set of dominant solutions that constitute the final Pareto front. That set is returned to the decision-maker, who can then select the most preferred one to be implemented.
Figure 4: The Pareto fronts of evolved prescriptors vs. heuristic baselines, with ELUC and change evaluated on the global test set. The trained prescriptors achieve better solutions than the baselines in the middle-change region where the changes matter the most, demonstrating that they can take advantage of nonlinear relationships in land use to discover useful, non-obvious solutions.
To demonstrate the value of evolutionary search, the prescriptors in the Pareto front were compared against two heuristic baselines (Figure 4). Each of them converts as much land as possible to secdf, which has the lowest ELUC, up to a set threshold. The first is a simple “even” heuristic, where an equal proportion of land equal to the threshold is taken from all land types and converted to secdf. The second is the optimal linear heuristic, where as much land as possible is first taken from the land type that has the highest weight in the global LinReg model, then from the land type that has the second highest weight, and so on until the given change threshold is reached. Ten such heuristic prescriptors were created with evenly distributed change thresholds in order to generate heuristic Pareto fronts. All three methods were then evaluated on a random 1% sample of the global test set from [2012-2021].
As Figure 3 shows, all prescriptions converged in the high-change region of the Pareto front, where 100% forest prescriptions dominate. The heuristics performed better in the low-change region, where the predictors may not be accurate enough to distinguish between small changes. However, in the most useful middle-change region, the evolved prescriptors outperformed both heuristics, thus demonstrating that they are exploiting the nonlinear relationships well.
5.3 Interactive Evaluation
An interactive demo of the trained models is provided at https://landuse.evolution.ml. It allows the user to explore different locations, observe the prescribed actions for them, and see their outcomes. It is also possible to modify those actions and see their effect, thus evaluating the user’s skills as a decision-maker compared to that of machine learning.
As an example, Figure 5 shows the grid cells from which the user has selected one. The system’s recommends a change in land-use that results in a 26.5% reduction in carbon emissions with a 22.0% change. Using the sliders, the user can then explore alternative tradeoffs and modify specific elements of each suggestion. For each variation, the user can then obtain the estimated carbon reduction and change percentage, and in this manner gain an understanding of the possible attainable solutions and confidence in the preferred choice.
Figure 5: A suggested land-use change for a given location. ($a$) The location is indicated by the red dot among the UK grid cells. ($b$) One prescriptor is chosen from the Pareto front spanning minimal change and minimal ELUC (top left). The current land use is shown on the left pie chart and the recommended one on the right pie chart. This prescriptor recommends decreasing pasture and annual C3 crops and increasing secondary forest and secondary non-forest vegetation, resulting in a 26.5% decrease in carbon emissions with a 22.0% change. The user can then select different solutions from the Pareto front and modify the sliders manually to explore alternatives.
6 Future Work
The most immediate future work consists of improving the predictor accuracy through ensembling. Given that models trained on specific regions perform the best on those regions, there is a good chance that an ensemble of them will do even better. Moreover, there may be benefit of including LinReg, RF, and neural network versions of these models in the ensemble as well. There are several ensembling techniques in the Project Resilience codebase, such as averaging, sampling, and confidence-based ensembling, and contributors are encouraged to submit more in the future. The effect of such alternatives will be evaluated in future work, both on prediction accuracy and on driving the search for better prescriptors. Also, a metric will be developed for prescriptor performance, such as the hypervolume of the Pareto front, making it possible to compare different predictor and prescriptor approaches quantitatively.
The predictor creates point estimates of ELUC. A technique called Residual Input-Output kernel (RIO; Qiu et al., 2020) can be used to estimate confidence in these predictions. The method builds a Gaussian Process model of the residual error, quantifying the uncertainty. RIO is part of the Project Resilience codebase but was not yet used in the study in this paper.
Project Resilience also provides an alternative approach to prescription: representing the prescriptor as a set of rules (Shahrzad et al., 2022). This approach has the advantage that the behavior is transparent and explainable, as demonstrated in prior experiments in other tasks (including NPI optimization (Miikkulainen et al., 2021)). Rule-set evolution on land-use optimization will be evaluated in the future.
To make the system more immediately useful to decision-makers, the cost metric could be refined to take into account region-specific costs of changing each particular land type to another. The approach can also be extended with preferences and further objectives. For instance, the decision maker could choose which crops should be preferred, which may be more actionable than the large-scale changes in the current dataset. It may be possible to optimize additional objectives such as minimizing nitrate pollution and water consumption, and maintaining or increasing food production. Similarly, instead of planning for a single year, it might be possible to develop prescriptors that recommend land-use changes for the next several years: For instance, what could be done in the next five years to reduce ELUC by a given target percentage. As data and simulations become more accurate in the future, such refined and extended decisions should become possible.
7 Conclusion
Land-use policy is an area of climate change where local decision-makers can have a large impact. In this paper, historical data and simulation technology are brought together to build an efficient machine-learning model that can predict the outcomes of these decisions efficiently for different contexts. An evolutionary optimization process is then used to identify effective policies, balancing the cost of land-use change and its effect in reducing carbon emissions. Machine learning methods can thus play an important role in empowering decision-makers on climate change issues.
Acknowledgements
We’d like to thank the BLUE, LUH2, and Project Resilience teams, in particular Amir Banifatemi, Prem Krishnamurthy, Gyu Myoung Lee, Gillian Makamara, Michael O’Sullivan, Julia Pongratz, Clemens Schwingshackl, and Jennifer Stave.