### v5

# Discovering Effective Policies for Land-Use Planning with Neuroevolution

Abstract.
How areas of land are allocated for different uses, such as forests, urban areas, and agriculture, has a large effect on the terrestrial carbon balance, and therefore climate change. Based on available historical data on land-use changes and a simulation of the associated carbon emissions and removals, a surrogate model can be learned that makes it possible to evaluate the different options available to decision-makers efficiently. An evolutionary search process can then be used to discover effective land-use policies for specific locations. Such a system was built on the Project Resilience platform (ITU, 2023) and evaluated with the Land-Use Harmonization dataset LUH2 (Hurtt et al., 2020) and the bookkeeping model BLUE (Hansis et al., 2015). It generates Pareto fronts that trade off carbon impact and amount of land-use change customized to different locations, thus providing a potentially useful tool for land-use planning.
Key words and phrases:
Surrogate modeling, Decision making, Climate Change, Land-use Change, Neural Networks, Evolutionary Computation
{CCSXML}
<ccs2012> <concept> <concept_id>10010147.10010257.10010321</concept_id> <concept_desc>Computing methodologies Machine learning algorithms</concept_desc> <concept_significance>500</concept_significance> </concept> <concept> <concept_id>10010147.10010919.10010172</concept_id> <concept_desc>Computing methodologies Distributed algorithms</concept_desc> <concept_significance>100</concept_significance> </concept> <concept> <concept_id>10010147.10010257.10010293.10011809.10011812</concept_id> <concept_desc>Computing methodologies Genetic algorithms</concept_desc> <concept_significance>500</concept_significance> </concept> <concept> <concept_id>10010147.10010257.10010293.10010294</concept_id> <concept_desc>Computing methodologies Neural networks</concept_desc> <concept_significance>500</concept_significance> </concept> </ccs2012>
1. Introduction
One of the main factors contributing to climate change is CO2 emissions due to land-use change (Friedlingstein et al., 2023). Land-use CO2 emissions crucially depend on how much land area is allocated for different uses. Forests in general take up more carbon than e.g. croplands and rangelands, yet such uses are essential for the economy. Land-use patterns must therefore be planned to maximize carbon removals while maintaining economic viability.
An approach for land-use optimization was developed as part of Project Resilience, a non-profit project hosted by the ITU agency of the United Nations (ITU, 2023). The goal is to provide decision-makers with a tool to know how their land-use choices affect CO2 fluxes in the long term, and make suggestions for optimizing these choices.
More specifically, the tool is designed to answer three questions: (1) For a geographical grid cell, identified by its latitude and longitude, what changes to the land use can be made to reduce CO2 emissions? (2) What will be the long-term CO2 impact of changing land use in a particular way? (3) What are the optimal land-use choices that can be made with minimal cost and maximal effect?
The approach is based on the Evolutionary Surrogate-assisted Prescription method (Francon et al., 2020). The idea is to first utilize historical data to learn a surrogate model on how land-use decisions in different contexts affect carbon emissions, and then use this model to evaluate candidates in an evolutionary search process for good land-use change policies.
As a result, a Pareto front is generated of solutions that trade off reduction in carbon emissions and the amount of change in land use. Each point in the Pareto front represents an optimal policy for that tradeoff. To make the results trustworthy, the tool allows the decision-maker to explore modifications to these policies, and see the expected effect. In the future, it should also be possible to evaluate confidence of the predictions, evolve rule sets to make the policies explainable, and utilize ensembling and further objectives and preferences to make them more accurate. Thus, the tool harnesses several techniques in machine learning to provide a practical tool for decision-makers in optimizing land-use decisions.
Project code and data are available at https://github.com/Project-Resilience/mvp/tree/main/use_cases/eluc, and an interactive demo at https://landuse.evolution.ml.
2. Background
The surrogate-assisted evolutionary optimization method used on the project is first outlined below, followed by the project background as the first application of Project Resilience.
2.1. Evolutionary Surrogate-assisted Prescription
Figure 1. The Evolutionary Surrogate-assisted Prescription (ESP) method for decision optimization. A predictor is trained with historical data on how given actions in given contexts led to specific outcomes. It is then used as a surrogate in order to evolve prescriptors, i.e. neural networks that implement decision policies resulting in the best possible outcomes.
Evolutionary Surrogate-assisted Prescription (ESP; Francon et al., 2020) is an approach for optimizing decision-making in a variety of domains (Figure 1). The main idea is that a decision policy can be represented as a neural network, or a set of rules, and a good policy can be discovered through population-based search, i.e. using evolutionary computation techniques. However, each candidate must be evaluated, which is difficult to do in many real-world applications. Therefore, a surrogate model of the world is learned from historical data, predicting how good the resulting outcomes are for each decision in each context.
More formally, given a set of possible contexts $\mathbb{C}$ and possible actions $\mathbb{A}$, a decision policy $D$ returns a set of actions $A$ to be performed in each context $C$:
$D(C)=A\;,$ (1)
where $C\in\mathbb{C}$ and $A\in\mathbb{A}$. For each such $(C,A)$ pair there is a set of outcomes $O(C,A)$, and the Predictor $P_{d}$ is defined as
$P_{d}(C,A)=O,$ (2)
and the Prescriptor $P_{s}$ implements the decision policy as
$P_{s}(C)=A\;,$ (3)
such that $\sum_{i,j}O_{j}(C_{i},A_{i})$ over all possible contexts $i$ and outcome dimensions $j$ is maximized (assuming increased outcomes are better). It thus approximates the optimal decision policy for the problem. The predictor can be learned from historical data on $CAO$ triples. In contrast, the optimal actions $A$ for each context $C$ are not known, and must therefore be found through search.
ESP was first evaluated in reinforcement learning tasks such as function approximation, cart-pole control, and the flappybird game, and found to discover significantly better solutions, find them faster, and with lower regret than standard approaches such as direct evolution, Deep Q-Networks (DQN; Mnih et al., 2015), and Proximal Policy Optimization (PPO; Schulman et al., 2017). Its power comes from automatic regularization and automatic curricular learning: When predictors are learned at the same time as the prescriptors, each prescriptor is evaluated against multiple predictors; as the predictors improve, they provide more refined evaluations. ESP is thus a powerful mechanism for learning policies in time-varying domains.
2.2. Project Resilience
A major application of ESP to decision-making was developed for optimizing strategies for non-pharmaceutical interventions (NPIs) in the COVID-19 pandemic (Miikkulainen et al., 2021). Using case and NPI data from around the world, collected in 2020-2022 by Oxford University’s COVID-19 project (Hale et al., 2020), a neural network model was developed to predict the course of the pandemic; this model was then used as a surrogate in evolving another neural network to prescribe NPIs. It discovered different strategies at different phases of the pandemic, such as focusing on schools and workplaces early on, alternating policies over time, and focusing on public information and masking later. This application demonstrated the power of ESP in discovering good tradeoffs for complex real-world decision-making tasks.
Encouraged by this application, the XPRIZE Pandemic Response Challenge (XPRIZE, 2023) was developed in 2020-2021 for NPI optimization using the same predictor+prescriptor approach. The competitors used several different machine learning or other techniques, and drew significant attention to AI-supported decision-making. The results were effective and in some cases used to inform decision-makers (e.g. on the 2021 surge in Valencia, Spain, and on the fall 2021 school openings in Iceland). An important conclusion was that it was possible to convene a group of experts—data scientists, epidemiologists, and public health officials—to build a useful set of tools to advise decision-makers on how to cope with and plan around a health disruption to society.
This experience with the XPRIZE Pandemic Response Challenge then led to Project Resilience (ITU, 2023). The XPRIZE competition serves as a blueprint for a collaborative effort to tackle global problems. The goal of Project Resilience is to build a public AI utility where a global community of innovators and thought leaders can enhance and utilize a collection of data and AI approaches to help with better preparedness, intervention, and response to environmental, health, information, social equity, and similarly scoped problems. The project is hosted by the ITU agency of the United Nations, under its Global Initiative for AI and Data Commons, and contributes to its general efforts toward meeting the Sustainable Development Goals (SDGs).
Staffed by volunteers from around the world, Project Resilience was started in 2022 with the design of a minimum viable product (MVP) platform. The goal is to
  • Develop an architecture to pull input and output data hosted by third parties;
  • Build an application programming interface (API) for third parties to submit models.
  • Develop code to compare both predictors and prescriptors in third-party models and produce a set of performance metrics;
  • Build a portal to visualize the assessment of predictors and prescriptors to include generations of key performance indicators (KPIs) and comparison across models;
  • Develop an ensemble model for predictors and prescriptors; and
  • Provide a user interface to decision-makers to help them in their decision-making process.
Land-use optimization is the first application of this MVP, as described below.
3. Land-Use Optimization Task
The data sources are described first, followed by how they are used as the context, action, and outcome variables. The elements of the approach include the predictor and prescriptor models, a method for modeling uncertainty, and ensembling the results.
3.1. Data
The data for carbon emissions (Emissions resulting from Land-Use Change, ELUC) originate from the Bookkeeping of Land-Use Emissions model (BLUE; Hansis et al., 2015), which attributes carbon fluxes to land-use activities. BLUE is one of three bookkeeping models that are used to provide ELUC estimates for the annually published Global Carbon Budget (GCB; Friedlingstein et al., 2023). Out of these three models, BLUE is the only one that provides spatially explicit ELUC estimates. In this paper, BLUE is used to estimate committed emissions due to land-use changes, which means that all the emissions or removals caused by a land-use change event are attributed to the year of the event. The emission (or removal) estimates thus relate to the long-term impact of a land-use change event, i.e. up to multiple decades (Hansis et al., 2015; Houghton et al., 1983). While in principle BLUE can be used as the surrogate model for ESP, in practice its calculations are too expensive to carry out on demand during the search for good policies. Therefore, we performed a number of BLUE simulations covering a comprehensive set of land-use changes for 1850-2022, resulting in a dataset that could be used to train an efficient surrogate model.
The Land-Use Change (LUC) data is provided by the Land-Use Harmonization project (LUH2; Hurtt et al., 2020). LUH2 provides data on fractional land-use patterns, underlying land-use transitions, and key agricultural management information, annually for the time period 850-2100 at 0.25$\times$0.25 degree resolution. LUH2 provides separate layers for states, transitions, and management. Version LUH2-GCB2022 from the Global Carbon Budget 2022 (Friedlingstein et al., 2022) was the basis for this paper. It is based on the HYDE3.3 historical database Klein Goldewijk et al. (2017); Kees Klein Goldewijk and van Zanden (2017) and uses the 2021 wood harvest data provided by the FAO agency of the United Nations. It consists of the following land-use types:
Primary: Vegetation that is untouched by humans
    primf: Primary forest
    primn: Primary nonforest vegetation
Secondary: Vegetation that has been touched by humans
    secdf: Secondary forest
    secdn: Secondary nonforest vegetation
Urban
    urban: Urban areas
Crop
    crop: All types pooled into one category
Pasture
    pastr: Managed pasture land
    range: Grazed natural land (i.e., grazing on grassland, savannah, etc.).
BLUE calculates yearly ELUC estimates based on the land-use changes provided by LUH2. Urban areas are assumed to have the same carbon content as croplands. All crop types are currently pooled into a single category. BLUE considers that the degradation of primary to secondary land often lowers the carbon content in vegetation and soils, and it considers pastures and the degradation of natural grasslands and savannahs by grazing. A complete overview of the model can be found in Hansis et al. (2015).
Together the ELUC and LUC datasets form the basis for constructing the context, action, and outcome variables, as will be described next.
3.2. Decision-making Problem
The modeling approach aims to understand the domain in two ways: (1) In a particular situation, what are the outcomes of the decision maker’s actions? (2) What are the actions that result in the best outcomes, i.e. the lowest carbon emission (or highest removals) with the smallest change in land use? The data is thus organized into context, action, and outcome variables.
Context describes the problem the decision maker is facing, i.e. a particular grid cell, a point in time when the decision has to be made, and the current land use at that time. More specifically it consists of:
  • Latitude and Longitude, representing the cell on the grid.
  • Area, representing the surface of the cell. Cells close to the equator have a bigger area than cells close to the poles.
  • Year, useful to capture historical decisions: The same cell may have been through many land-use changes over the years.
  • Land use, representing the percentage of land in the cell used by each type above. For simplicity, all crop types were pooled into a single type. In addition, a ‘nonland’ type represents the percentage of the cell that is not land (e.g. typically sea, lake, etc.). The percentages sum up to 100% in each grid cell.
Actions represent the choices the decision-maker faces. How can they change the land? In the study of this paper, these decisions are limited in two ways:
First, decision-makers cannot affect primary land. The idea is that it is always better to preserve primary vegetation, such as pristine rainforest; destroying it is not an option given to the system. Technically, it is not possible to re-plant primary vegetation. Once destroyed, it is destroyed forever. If re-planted, it would become secondary vegetation, for which BLUE assumes lower carbon content than for primary vegetation Hansis et al. (2015).
Second, decision-makers cannot affect urban areas. The needs of urban areas are dictated by other imperatives, and optimized by other decision makers. Therefore, the ESP system cannot recommend that a city should be destroyed, or expanded.
Outcomes consist of two conflicting variables. The primary variable is ELUC, i.e. emissions from land-use change. It consists of all CO2 emissions attributed to the land-use change, in metric tons of carbon per hectare (tC/ha), obtained from the BLUE simulation. Note that all follow-up CO2 emissions (or removals) are attributed to the year of the land-use change, thus neglecting temporal dynamics in CO2 fluxes. A positive number means carbon is emitted to the atmosphere and a negative number means carbon is taken up by vegetation. The secondary variable is the cost of the land-use change, represented by the percentage of land that was changed. This variable is calculated directly from the actions.
There is a tradeoff between these two objectives: It is easy to reduce emissions by changing most of the land, but that would come at a huge cost. Therefore, decision-makers have to minimize ELUC while minimizing land-use change at the same time. Consequently, the result is not a single recommendation, but a Pareto front where each point represents the best implementation of each tradeoff given a balance between the two outcomes.
Model Time (s) EU SA US Global
LinReg (EU) 0.047 0.033 0.172 0.169 0.206
LinReg (SA) 0.457 0.137 0.153 0.061 0.110
LinReg (US) 0.331 0.139 0.146 0.035 0.073
LinReg (Global) 4.644 0.139 0.150 0.035 0.074
RF (EU) 17.697 $0.064$ $0.211^{{\dagger}}$ $0.161^{{\dagger}}$ $0.218^{{\dagger}}$
RF (SA) 209.688 $0.133^{*{\dagger}}$ $\textbf{0.071}^{*{\dagger}}$ $0.074^{{\dagger}}$ $0.126^{{\dagger}}$
RF (US) 111.701 $0.163$ $0.185^{{\dagger}}$ $0.032^{*}$ $0.094^{{\dagger}}$
RF (Global) 417.647 $0.041^{*{\dagger}}$ $0.076^{*{\dagger}}$ $0.028^{*}$ $\textbf{0.045}^{*{\dagger}}$
NeuralNet (EU) 10.711 $\textbf{0.025}^{*{\dagger}}$ $0.277$ $0.286$ $0.334$
NeuralNet (SA) 103.696 $0.248$ $0.100^{*}$ $0.562$ $0.399$
NeuralNet (US) 73.141 $0.136^{{\dagger}}$ $0.225$ $\textbf{0.024}^{*{\dagger}}$ $0.150$
NeuralNet (Global) 1649.193 $0.046^{*}$ $0.110^{*}$ $0.025^{*{\dagger}}$ $0.050^{*}$
Bold: the best model in each region; *: 99% confidence the model outperforms LinReg; †: 99% confidence that RF outperforms NeuralNet, or vice-versa
Table 1. Average training time and mean absolute error in tC/ha of the trained models of each type trained on each region and evaluated on that region as well as all other regions. Statistical significance was estimated in each case with the unpaired $t$-test. The NeuralNet models are more accurate RF on EU and US, RF more accurate than NeuralNets on SA and globally; both are more accurate than LinReg in all regions (except RF on EU). Overall, the specialized models are more accurate than the Global models. However, RF models do not extrapolate well (as seen in Figure 2), which is necessary for evaluating prescriptor candidates. In contrast, NeuralNets do, and they are also otherwise reasonably accurate. Therefore, the Global NeuralNet model was used to evolve the prescriptors.
4. Models
The system consists of the predictor, trained with supervised learning on the historical data, and the prescriptor, evolved against the predictor.
Prediction: Given the context and actions that were performed, the predictive model estimates the outcomes. In this case, since the cost outcome (i.e. the percentage of changed land) can be calculated directly, only the ELUC is predicted by the model. That is, given the land use of a specific location, and the land-use changes that were made during a specific year, the model predicts the long-term CO2 emissions directly caused by these land-use changes.
Any predictive model can be used in this task, including linear regression, random forest, or a neural network; they will each be evaluated below. As usual in machine learning, each model is fit to the existing historical data and evaluated with left-out data.
Prescription: Given context, the prescriptive model suggests actions that optimize the outcomes. The model has to do this for all possible contexts, and therefore it represents an entire strategy for optimal land use. The strategy can be implemented in various ways, like sets of rules or neural networks. The approach in this paper is based on neural networks.
The optimal actions are not known, but the performance of each candidate strategy can be measured (using the predictive model), therefore the prescriptive model needs to be learned using search techniques. Standard reinforcement learning methods such as PPO and DQN are possible; the experiments in this paper use evolutionary optimization, i.e. conventional neuroevolution (Stanley et al., 2019). As in prior applications of ESP (Francon et al., 2020; Miikkulainen et al., 2021), the network has a fixed architecture of two fully connected layers; its weights are concatenated into a vector and evolved through crossover and mutation.
5. Experiments
This section describes the main results of the project so far, including different ways to create good predictors and the evolution of effective prescriptors based on them. The behavior of the discovered prescriptors is outlined, and an interactive demo of the entire land-use recommendation system is described.
5.1. Prediction
In preliminary experiments, prediction performance was found to differ between major geographical regions. To make these differences explicit, separate models were trained on different subsets of countries: Western Europe (EU), South America (SA), and the United States (US), making up 1.12, 9.99, and 7.26 percent of the global dataset, respectively. The EU subset contains the UK, France, Germany, the Netherlands, Belgium, Switzerland, and Ireland. The SA subset includes Brazil, Bolivia, Paraguay, Peru, Ecuador, Colombia, Venezuela, Guyana, Suriname, Uruguay, Argentina, and Chile. These countries were chosen as a representative subset to make the experiments computationally less expensive.
Upon examining the global dataset, a linear relationship was found between ||||ELUC|||| and land-use change with an R-squared value of 0.92. Therefore, a linear regression (LinReg) was fit to the data as the first predictive model. In preliminary experiments, the LinReg model was found to perform the best when current land use, latitude, longitude, cell area, and time were removed from the inputs. Thus, the difference in land use was used as the only input feature in the experiments. Data from the years 1851–2011 (inclusive), totaling 38,967,635 samples globally, was used to train the LinReg models; the 2,420,350 remaining samples from 2012–2021 (inclusive) were used for testing.
Second, a random forest model (RF) was trained on the same data but now including the full set of input features, in order to take advantage of RF’s ability to decide which features are most useful. The forest consisted of 100 decision trees with unrestricted maximum depth; at each split, a random subset of $n$ features was considered (where $n$ is the square root of the total number of features). To make the training time of the Global RF model feasible, it was trained with data from only 1982–2011. Preliminary experiments showed that constraining the training set in this manner did not decrease performance significantly. Both the linear regression and random forest models were trained using the scikit-learn library Pedregosa et al. (2011).
Third, a neural network (NeuralNet) was trained for the same task with the full set of features and data from 1851–2011 using the PyTorch library Paszke et al. (2019). The network was fully connected, composed of an input layer, a single hidden layer of size 4096, and an output layer. The output layer took as input the concatenation of the input layer and the hidden layer in order to take advantage of the observed linear relationships more easily (Cheng et al., 2016). It was trained for three epochs with batch size 2048 using the AdamW optimization algorithm Loshchilov and Hutter (2019). After each epoch, the learning rate was scaled by a factor of 0.1.
($a$) LinReg predictions
($b$) RF predictions
($c$) NeuralNet predictions
Figure 2. Visualizing the differences in model behavior. Predictions are created for the Global models using synthetic data created by changing 100% of land-use type A (row) to 100% of type B (column) in a 1% sample across the range of latitude, longitude, cell area, and year occurring in the test data. The results are averaged for each conversion from A to B. The models generally agree on the sign of ELUC, suggesting that the results are reliable. The RF model is not able to extrapolate to large values, resulting in low predictions (shown in a much smaller color scale to still make them visible); LinReg and NeuralNet are similar but differ numerically, presumably due to the differences between linear and nonlinear predictions.
LinReg, RF, and NeuralNet models were trained separately on each region and globally. While LinReg is deterministic and needed to be trained only once, RF and NeuralNet are stochastic in their initialization and training and were therefore trained 30 times to evaluate statistical significance. All models were tested on all regions with data from the most recent years, i.e. 2012–2021. As shown in Table 1, the LinReg models were fast to train, but they performed the worst (i.e., they have the largest mean absolute error in almost all cases), suggesting that the linear relationship is not enough to account for the data fully. RF performed significantly better than LinReg in all regions but EU, significantly worse than NeuralNet in EU and US, and outperformed NeuralNets in SA and globally. However, RF models do not generalize/extrapolate well to large changes in land use that occur in the dataset infrequently (or not at all for some land types; Figure 2), which turns out to be important when evolving prescriptors. Many prescriptor candidates suggest large land-use changes, and the predictor model needs to be able to evaluate them. In contrast, NeuralNets capture nonlinearities, which explains why they outperform LinReg in all regions. Most importantly, they extrapolate well, which suggests that they are a better choice for the surrogate model in prescriptor evolution. Additionally, LinReg and NeuralNet models have many fewer parameters than RF (taking 4KB and 363KB on disk respectively vs. 14GB for RF).
In order to visualize the behavior of the predictor models, they were queried with a synthetic dataset created with every possible conversion from 100% land type A to 100% land type B (the types primf, primn, and urban were excluded from B as was explained in Section 3.2). Because the predictor takes latitude, longitude, cell area, and year as input, 24,204 (1%) latitude/longitude/cell-area/year combinations from the test dataset were sampled randomly and evaluated with each conversion. The results were averaged across the samples to get the overall result for the conversion. Figure 2 shows these results for the Global models. The first finding is that all models generally rank conversions to crops $<$ pasture $<$ forest in terms of reducing ELUC, and generally agree on the sign of ELUC. This consistency suggests that the results are reliable. Second, while LinReg and NeuralNet make similar predictions overall, their numerical values differ in many places, presumably due to the nonlinearities. Third, despite RF’s generally high accuracy in known scenarios, it is not able to generalize/extrapolate to large changes and instead predicts low-magnitude ELUC values for all imposed land-use transitions.
Thus, since NeuralNet achieves similar or better performance than the RF in most cases, extrapolates well to extremes, and can learn nonlinear relationships in the data, the Global NeuralNet was used as the surrogate model to evolve the prescriptors.
5.2. Prescriptions
The prescriptors were fully connected neural networks with two layers of weights, implemented with Keras/Tensorflow. The topology was fixed and only the weights were evolved. The input consisted of 12 float values, i.e. the eight land-use percentages and the latitude, longitude, and area of the geographical cell, as well as the year. The hidden layer contained 16 units with tanh activation.
($a$) Evolution of the Pareto front
($b$) Performance of all evaluated prescriptors
Figure 3. Evolution of prescriptors with the Global NeuralNet predictor. ($a$) The Pareto front moves towards the lower left corner over evolution, finding better implementations for the different tradeoffs of the ELUC and land-use change objectives. ($b$) Each prescriptor evaluated during evolution is shown as a dot, demonstrating a wide variety of solutions and tradeoffs. The final Pareto front (collected from all generations) is shown as red dots, constituting a set of solutions from which the decision-maker can choose a preferred one.
Figure 4. The Pareto fronts of Evolved Prescriptors vs. heuristic baselines, with ELUC and land-use change evaluated on the Global test set. The Evolved Prescriptors achieve better solutions than the baselines in the middle-change region where the land-use changes matter the most, demonstrating that they can take advantage of nonlinear relationships in land use to discover useful, non-obvious solutions.
($a$) Average performance
($b$) Performance on underlying samples
($c$) Performance differences
Figure 5. Comparing Evolved Prescriptors with the heuristics. ($a$) The average performance of the Evolved Prescriptor dominates those of the heuristics. ($b$) The averages are expanded into actual samples in the test set (subsampled for readability). The Evolved Prescriptor suggests many more large changes than the heuristics. ($c$) The differences in change percentage and ELUC between the Evolved Prescriptor and the Perfect Heuristic for each test sample, with color indicating which solution dominates. Surprisingly, the Evolved Prescriptor dominates the Perfect Heuristic only on a few individual samples. Thus, evolution discovered the insight that it is possible to do well globally by utilizing a few cases where large change is possible.
Figure 6. Characterizing the recommendations. The strategies used by the Evolved Prescriptors and the heuristics are illustrated by plotting the average amount of change in each land-use type. Primarily, they all suggest converting cropland to secondary forest. The Evolved Prescriptor removes more crop on average, and generally takes advantage of more change.
During evolution, prescriptor candidates were evaluated on a random subset of 1% of the Global dataset from years 1851–2011 with the Global NeuralNet predictor. Prescriptor output was a single vector of size five (the eight land-use percentages minus primf, primn, and urban, which were considered fixed), produced by the ReLU activation function. These five outputs were normalized to sum to the previous amount of land taken by these five types. The result is the prescriptor’s suggestion for the percentage of land use of each type. The five outputs were combined with the fixed land-use values for primf, primn, and urban to compute the difference from the initial land use. This difference was concatenated with the initial land use as well as latitude, longitude, cell area, and year, and passed to the Predictor model to get the predicted ELUC metric. This metric was aggregated across the evaluation data set to obtain the performance outcome of the suggestion. The cost outcome, i.e. the percentage of land-use change, was computed based on the difference between the five outputs and the corresponding inputs. All positive differences were summed for a given cell and divided by the total land-use percentage of the cell (excluding nonland).
The aggregated ELUC and the land-use change outcomes were minimized with an evolutionary search on the neural network weights. Conventional neuroevolution with crossover and mutation operators on a vector of network weights was combined with NSGA-II (Deb et al., 2000) to account for multiobjectivity. To speed up evolution, two prescriptor models were trained with backpropagation and injected into the initial population: One to prescribe no land-use changes at all and the other to prescribe as much secondary forest (secdf) as possible (which was observed to reduce ELUC the most). The rest of the population was created randomly through an orthogonal initialization of weights in each layer with a mean of zero and a standard deviation of one (Saxe et al., 2014). Evolution was run for 100 generations with the following parameters:
  • nb_elites: 10
  • mutation_type: gaussian_noise_percentage
  • mutation_factor: 0.2
  • population_size: 100
  • parent_selection: tournament
  • initialization_range: 1
  • mutation_probability: 0.2
  • remove_population_pct: 0.8
  • initialization_distribution: orthogonal.
Figure 3 demonstrates the progress of evolution towards increasingly better prescriptors, i.e. those that represent better implementations of each tradeoff of the ELUC and land-use change objectives. These solutions will be referred to as "Evolved Prescriptors" in the rest of the paper. They represent a wide variety of tradeoffs and a clear set of dominant solutions that constitute the final Pareto front. That set is returned to the decision-maker, who can then select the most preferred one to implement.
To demonstrate the value of evolutionary search, the Evolved Prescriptors were compared against two heuristic baselines (Figure 4). Each of them converts as much land as possible to secdf, which has the lowest ELUC, up to a set threshold. The first is a simple baseline called "Even Heuristic", where an equal proportion of land up to the threshold is taken from all land-use types and converted to secdf. The second is an optimal linear baseline called "Perfect Heuristic", where as much land as possible is first taken from the land-use type that has the highest weight in the Global LinReg model, then from the land-use type that has the second highest weight, and so on until either the given land-use change threshold is reached or there is no land left to convert to secdf. One hundred such heuristic solutions were created with evenly distributed land-use change thresholds between 0% and 100% in order to generate heuristic Pareto fronts. All three methods were then evaluated on a random 1% sample of the Global test set from the years 2012–2021.
As Figure 3 shows, the Evolved Prescriptors and the heuristics all converged in the high-change region of the Pareto front, where 100% forest prescriptions dominate. The heuristics performed slightly better in the low-change region, where the predictors may not be accurate enough to distinguish between small changes. However, in the most useful middle-change region, the Evolved Prescriptors outperformed both heuristics, thus demonstrating that they are likely exploiting the nonlinear relationships well. The performance for each prescription method can be quantified in terms of hypervolume, which is measured as the rectangular area above and to the right of each point (only counting overlapping area once), extending to ELUC = 0 and change = 1. The Even Heuristic is outperformed by the Perfect Heuristic which is outperformed by the Evolved Prescriptors, with hypervolumes of 19.68, 20.30, and 20.52, respectively.
5.3. Prescriptor Behaviors
[…]
($a$) The location of the cell
($b$) Suggestions of one Evolved Prescriptor chosen from the Pareto front
Figure 7. A suggested land-use change for a given location (screenshots from the demo at https://landuse.evolution.ml). ($a$) The location is indicated by the red dot among the UK grid cells. ($b$) One Evolved Prescriptor is chosen from the Pareto front spanning minimal change and minimal ELUC (top left). The current land use is shown on the left pie chart and the recommended one on the right pie chart, as well as the sliders on the left. This prescriptor recommends decreasing pasture and crops and increasing secondary forest, resulting in an 18.09 tC/ha decrease in carbon emissions with a 15.77% land-use change. The user can then select different solutions from the Pareto front and modify the sliders manually to explore alternatives.
[…]
The reason is that evolution discovered a way to take advantage of the large changes. In order to perform better on average, it does not need to outperform the Perfect Heuristic at the individual cell level; instead, it can identify and utilize cases where large changes make a big difference. Thus, even though the Evolved Prescriptor almost never beats the Perfect Heuristic in any individual cell, when averaged across all cells, it dominates the heuristic. In other words, evolution discovered the insight that it is most important to focus land-use changes in those regions where the largest changes are possible. Pick your battles right, and you’ll win overall.
In order to see what the recommendations look like, the average policies of the three approaches are illustrated in Figure 6. The greatest amount of negative change made is to crop and positive is to secondary forest, i.e. all approaches recommend converting cropland to forest, which makes sense.
5.4. Interactive Evaluation
An interactive demo of the trained/evolved models is provided at https://landuse.evolution.ml. It allows the user to explore different locations, observe the prescribed actions for them, and see their outcomes. It is also possible to modify those actions and see their effect, thus evaluating the user’s skills as a decision-maker compared to that of machine learning.
As an example, Figure 7 shows the grid cells from which the user has selected one. The system recommends a change in land use that results in an 18.09 tC/ha reduction in carbon emissions with a 15.77% land-use change. Using the sliders, the user can then explore alternative tradeoffs and modify specific elements of each suggestion. For each variation, the user can then obtain the estimated carbon reduction and land-use change percentage, and in this manner gain an understanding of the possible attainable solutions and confidence in the preferred choice.
6. Future Work
In the project so far we have not distinguished between the different types of crops and only partly considered regional variations in how land-use changes affect carbon stocks in soil and vegetation (by training different models for different regions). The assumed values for carbon stocks in soil and vegetation represent conditions of the recent past (Hansis et al., 2015); they do not consider environmental effects such as disturbances due to fires or droughts or better vegetation growth due to CO2 effects or a warmer climate, which can substantially impact terrestrial CO2 fluxes (e.g. Aragão et al., 2018; Pongratz et al., 2021; Rosan et al., 2024). Further, estimates of CO2 fluxes from land-use change also depend on the spatial resolution of the land-use change dataset (Ganzenmüller et al., 2022; Bultan et al., 2022). Extending the project with such details should lead to more precise and more fine-scale estimates of the impact of land-use change decisions in future work.
Another immediate future work direction consists of improving the predictor accuracy through ensembling. Given that models trained on specific regions perform the best on those regions, there is a good chance that an ensemble of them will do even better. Moreover, there may be a benefit of including LinReg, RF, and NeuralNet versions of these models in the ensemble as well. There are several ensembling techniques in the Project Resilience codebase, such as averaging, sampling, and confidence-based ensembling, and contributors are encouraged to submit more in the future. The effect of such alternatives will be evaluated in future work, both on prediction accuracy and on driving the search for better prescriptors. The hypervolume of the Pareto front can be used to compare the different approaches quantitatively.
The predictor creates point estimates of ELUC. A technique called Residual Input-Output kernel (RIO; Qiu et al., 2020) can be used to estimate confidence in these predictions. The method builds a Gaussian Process model of the residual error, quantifying the uncertainty. RIO is part of the Project Resilience codebase but was not yet used in the study in this paper.
Project Resilience also provides an alternative approach to prescription: Representing the prescriptor as a set of rules (Shahrzad et al., 2022). This approach has the advantage that the behavior is transparent and explainable, as demonstrated in prior experiments in other tasks (including NPI optimization (Miikkulainen et al., 2021)). Rule-set evolution on land-use optimization will be evaluated in the future.
To make the system more immediately useful to decision-makers, the cost metric could be refined to take into account region-specific costs of changing each particular land-use type to another. The approach can also be extended with preferences and further objectives. For instance, the decision maker could choose which crops should be preferred, which may be more actionable than the large-scale changes in the current dataset. It may be possible to optimize additional objectives such as minimizing nitrate pollution and water consumption, and maintaining or increasing food production. Similarly, instead of planning for a single year, it might be possible to develop prescriptors that recommend land-use changes for the next several years: For instance, what could be done in the next five years to reduce ELUC by a given target percentage. As data and simulations become more accurate in the future, such refined and extended decisions should become possible.
7. Conclusion
Land-use policy is an area relevant to climate change where local decision-makers can have a large impact. In this paper, historical data and simulation technology are brought together to build an efficient machine-learning model that can predict the outcomes of these decisions efficiently for different contexts. An evolutionary optimization process is then used to identify effective policies, balancing the cost of land-use change and its effect in reducing carbon emissions. Machine learning methods can thus play an important role in empowering decision-makers to act on climate change issues.
Acknowledgements
We would like to thank the BLUE, LUH2, and Project Resilience teams, in particular Amir Banifatemi, Prem Krishnamurthy, Gyu Myoung Lee, Gillian Makamara, Michael O’Sullivan, Julia Pongratz, and Jennifer Stave.